OpenStudy (anonymous):

I need these asap What are the classifications of each system? 1. x + 5y = - 2 x + 5y = 4 2. y = 3x + 4 - 2x + y = 4 3. y = - x + 3 2x + 2y = 6 4. 5x + 2y = - 4 5x - 2y = 3 A. consistent independent B. coincident C. inconsistent

OpenStudy (sloppycanada):

Well, one at a time - Inconsistent. Has no solution. Let me go through and solve each of these.

OpenStudy (sloppycanada):

A = inconsistent.

OpenStudy (anonymous):

thank you for helping me

OpenStudy (anonymous):

did you mean number one is inconsistent you put a

OpenStudy (sloppycanada):

Sorry yeah.

OpenStudy (anonymous):

k just take your time solving them im in no rush i just suck at math lol

OpenStudy (sloppycanada):

Wait, you have four problems and only three types of problems.

OpenStudy (anonymous):

you can use the abc more than once

OpenStudy (anonymous):

you can use the answer choices more than once

OpenStudy (sloppycanada):

I have never heard of that first one, consistent independent. I'm sorry, not sure i can help you out.

OpenStudy (anonymous):

@FaireGaming

OpenStudy (zzr0ck3r):

A means every ordered pair (x,y) is a solution B means there is one solution C means there are no solutions

OpenStudy (zzr0ck3r):

Do you know how to solve systems of linear equations?

OpenStudy (anonymous):

i already knew that my brain wont let me figure out which one is which and yes

OpenStudy (anonymous):

@FaireGaming

OpenStudy (zzr0ck3r):

that is all you need to do. What do you get when you solve the first one?

OpenStudy (anonymous):

give me one sec to solve it

OpenStudy (anonymous):

i take that back i dont know how to solve it im completely confused

OpenStudy (zzr0ck3r):

ok lets do the first one

OpenStudy (anonymous):

i really dont mean to be rude but its 1 am and i have to go to sleep soon

OpenStudy (zzr0ck3r):

x+5y=-2 Solve this for x x=-2-5y Now plug this in to the second equation for x x+5y=4 becomes -2-5y+5y=4 and we get -2=4 Since this makes no sense, it is inconsistent

OpenStudy (zzr0ck3r):

that is not rude....

OpenStudy (anonymous):

im so confused that makes no sense

OpenStudy (zzr0ck3r):

so if you get something like -2=4 it is inconsistent if you get something like 2=2 it is consistent if you get something like y=3 or x=5 then it is linearly consistent

OpenStudy (zzr0ck3r):

which part?

OpenStudy (anonymous):

i think i got it

OpenStudy (zzr0ck3r):

try the second one

OpenStudy (zzr0ck3r):

notice that the first equation for part 2 is already solved for y, so just plug it in for y in the second equation

OpenStudy (zzr0ck3r):

simplify and tell me what you get

OpenStudy (anonymous):

so the answer is 1a 2b 3b 4 c does that sound right thats what i got when i tried it

OpenStudy (anonymous):

does that sound right?

OpenStudy (zzr0ck3r):

no, and I find it very hard to believe you did all 4 systems in one minute

OpenStudy (zzr0ck3r):

I teach this class and I could not...

OpenStudy (zzr0ck3r):

If you take the time to do what I am showing you will spend much less time then trying to get me to give you answers. Sorry if that sounded rude :)

OpenStudy (anonymous):

im really sorry its just im really tired and i have dyslexia so my head mixes up all the numbers so math is really hard for me

OpenStudy (anonymous):

1 is c right did i do that right or am i still wrong i really am trying

OpenStudy (anonymous):

@zzr0ck3r

OpenStudy (zzr0ck3r):

correct

OpenStudy (zzr0ck3r):

for the second one take the first equation that says y=-x+3 and plug that into the equation below it for y, in other words replace y in the second equation with -x+3 tell me what you have after you do that

OpenStudy (anonymous):

those numbers are for the 3rd one i thought we where doing the second one

OpenStudy (anonymous):

oh crap its 130 am im getting in trouble because im not asleep i gotta go this sucks i didnt finish @zzr0ck3r

OpenStudy (zzr0ck3r):

do the same things for the second one

OpenStudy (anonymous):

i cant i have to go to sleep my moms yelling at me

OpenStudy (zzr0ck3r):

ok good night. stop by tomorrow and we can complete it

OpenStudy (liv1234):

@ashley011802 Maybe you shouldn't wait to do your school work until 1 am, no offense.

OpenStudy (zzr0ck3r):

well, it could be the case that she had been working on it all night.

OpenStudy (zzr0ck3r):

I hope :)

OpenStudy (anonymous):

thats the problem it was due tonight i do online school now im gonna fail it because i have to go to sleep and i was doing it al day @Liv1234 stay out of it

OpenStudy (liv1234):

True, but still.

OpenStudy (zzr0ck3r):

since you said you had to go to sleep you could have solved this 3 times

OpenStudy (liv1234):

Please don't be rude, I was just simply stating, I'm already dealing with the death of a friend from yesterday, I do online school too and I know it can be stressful.

OpenStudy (anonymous):

no i cant i metaly cant with dyslexia

OpenStudy (zzr0ck3r):

I have dyslexia and I teach math at university

OpenStudy (anonymous):

@zzr0ck3r your being really rude to me im not you

OpenStudy (zzr0ck3r):

how is that rude?

OpenStudy (zzr0ck3r):

I am saying that you have spent at least 20 minutes now saying that you dont have time to do the problem

OpenStudy (liv1234):

The point is, is that you shouldn't wait this late to do your work. If you do Connections Academy, they don't fail you for having late work.

OpenStudy (zzr0ck3r):

lets not argue over old things

OpenStudy (zzr0ck3r):

math is why we are hear

OpenStudy (zzr0ck3r):

I really am dyslexic also, I had to take special classes growing up for it

OpenStudy (anonymous):

i dont i do a private online school and if its not turned in by the dead line they fail you and ive been doing school work all day i didnt wait

OpenStudy (zzr0ck3r):

do you understand that we could be doing the math right now instead of talking about this?

OpenStudy (anonymous):

hey sorry my mom said i really have to get off or shes taking my computer away

OpenStudy (zzr0ck3r):

ok then